# Computable Thermodynamic Formalism

### v1

@@ 6.3. Computable expanding Thurston maps
We begin by constructing a recursively enumerable set $K$ and a sequence $\{Y_{n,k}\}_{(n,k)\in L}$ of uniformly lower semi-computable open sets, where $L=\mathbb{N}\times K$. Now we write $\mathcal{S}\bigl(\widehat{\mathbb{C}}\bigr)=\{s_{i}\}_{i\in\mathbb{N}}$. Since $\operatorname{post}{F}$ is a finite set of computable points, by Proposition 3.20 (i), (ii), and (v), $\widehat{\mathbb{C}}\smallsetminus\operatorname{post}{F}$ is a nonempty lower semi-computable open set. By Proposition 3.8, there exists a recursively enumerable set $I\subseteq\mathbb{N}$ such that $\mathbb{Q}\bigl(\widehat{\mathbb{C}}\bigr)\smallsetminus\operatorname{post}{F}=\{s_{i}:i\in I\}$. We define $r_{i}\coloneqq\sigma(s_{i},\operatorname{post}{F})>0$ for each $i\in I$. Thus, by Definition 3.3, $\{r_{i}\}_{i\in I}$ is uniformly computable. Consequently, by Proposition 3.10, $\{D_{i}\}_{i\in I}$ is uniformly lower semi-computable open, where $D_{i}=B(s_{i},r_{i})$ for each $i\in I$. As $F$ is computable, $\bigl\{F^{-1}(D_{i})\bigr\}_{i\in I}$ is also uniformly lower semi-computable. By Lemma 6.9, there exists a recursively enumerable set $K\subseteq\mathbb{N}\times I$ and a sequence $\{V_{k}\}_{k\in K}$ of uniformly lower semi-computable open sets that are connected such that $\bigcup_{k\in K_{i}}V_{k}=F^{-1}(D_{i})$ for each $i\in I$.
@@
We argue by contradiction that for each pair of distinct integers $1\leqslant l,\,l^{\prime}\leqslant b$,
$h_{l}(D_{i})\cap h_{l^{\prime}}(D_{i})=\emptyset.$ (6.7)
We assume that (6.7) is not true for some pair of distinct integers $1\leqslant l,\,l^{\prime}\leqslant b$. Then there exist $z,\,z^{\prime}\in D_{i}$ with $h_{l}(z)=h_{l^{\prime}}(z^{\prime})$. Hence, $z=F(h_{l}(z))=F(h_{l^{\prime}}(z^{\prime}))=z^{\prime}$. Combined with [BM17, Lemma A.6 (i)], this implies that $h_{l}=h_{l^{\prime}}$, which contradicts $l\neq l^{\prime}$. Thus, (6.7) is true for each pair of distinct integers $1\leqslant l,\,l^{\prime}\leqslant b$.
@@
We can enumerate $\mathrm{Per}(F)\cap\operatorname{post}{F}$ by $\{y_{i}\}_{i=1}^{N}$ as a sequence of uniformly computable points that are different from each other, since it is a finite set of computable points.
We now construct a sequence $\{\mathbb{B}_{i}\}_{i=1}^{N}$ of uniformly lower semi-computable open sets satisfying that $\{y_{i}\}\subseteq\mathbb{B}_{i}\subseteq W^{1}(y_{i})$ for each integer $1\leqslant i\leqslant N$. Fix an arbitrary integer $1\leqslant i\leqslant N$. By Remark 6.10, $W^{1}(y_{i})$ is an open set containing the point $y_{i}$. Then there exists $R_{i}\in\mathbb{Q}^{+}$ with $B(y_{i},R_{i})\subseteq W^{1}(y_{i})$. Since $y_{i}$ is a computable point, by Proposition 3.10, $\mathbb{B}_{i}\coloneqq B(y_{i},R_{i})$ is a lower semi-computable open set with $\{y_{i}\}\subseteq\mathbb{B}_{i}\subseteq W^{1}(y_{i})$. Hence, by Definition 3.7, $\{\mathbb{B}_{i}\}_{i=1}^{N}$ is uniformly lower semi-computable open.
@@
Proof of Claim 2. Now we verify Assumptions (ii) to (v) in Subsection 5.3.
Assumption (ii): Since $\phi_{0}$ is the constant function, this assumption holds.
@@
Assumption (iv): Note that $\widehat{\mathbb{C}}$ is recursively compact in $(X,\,\rho,\,\mathcal{S})$ by Proposition 6.5. Hence, by definitions, it suffices to show that $\operatorname{card}\mathopen{\big(}F^{-1}(\operatorname{post}{F})\mathclose{\big)}<+\infty$ and $\mathcal{E}_{0}(F,\phi_{0})\cap\mathcal{P}\bigl(X;\widehat{\mathbb{C}}\smallsetminus F^{-1}(\operatorname{post}{F})\bigr)=\{\mu_{0}\}$. Moreover, by Theorem 6.3 and [BM17, Corollary 17.3], we have $\mathcal{E}(F,\phi_{0})=\mathcal{E}(f,\phi_{0})=\{\mu_{0}\}$. Then we show that $\mu_{0}(\operatorname{post}{F})=0$. We argue by contradiction and assume that $\mu_{0}(\operatorname{post}{F})>0$. Combined with the fact that $\operatorname{card}\mathopen{(}\operatorname{post}{F}\mathclose{)}<+\infty$, this implies that $\mu_{0}(\{l_{0}\})>0$ for some $l_{0}\in\operatorname{post}{F}$. On the other hand, by [BM17, Theorem 17.1], the measure $\mu_{0}$ is mixing. Since mixing implies ergodicity, it follows that $\mu_{0}$ is ergodic. Hence, by [LS24b, Lemma 6.3 (iii)], it follows from $\mu_{0}(\{l_{0}\})>0$ that $l_{0}$ is a periodic point of $F$ and $\mu_{0}=\frac{1}{m}\sum_{k=0}^{m-1}\delta_{F^{k}(l_{0})}$ for some $m\in\mathbb{N}$. Thus, it follows from (3.1) that $h_{\mu_{0}}(F)=0$. However, since $\mu_{0}\in\mathcal{E}(F,\phi_{0})$, $h_{\mu_{0}}(F)=h_{\mathrm{top}}(F)=\log(\deg{F})>0$. This contradiction shows that $\mu_{0}(\operatorname{post}{F})=0$. Hence, by $\mu_{0}\in\mathcal{M}\bigl(\widehat{\mathbb{C}},F\bigr)$, we have $\mu_{0}\bigl(F^{-1}(\operatorname{post}{F})\bigr)=\mu_{0}(\operatorname{post}{F})=0$. Thus, we obtain that $\mathcal{E}_{0}(F,\phi_{0})\cap\mathcal{P}\bigl(X;\widehat{\mathbb{C}}\smallsetminus F^{-1}(\operatorname{post}{F})\bigr)=\{\mu_{0}\}$.
@@
Assumption (iii): Since $F$ is an expanding Thurston map, by [BM17, Corollary 17.2], we have $h_{\mathrm{top}}(F)=\log(\deg{F})$. Suppose $h=\phi_{0}$. Then property (i) in Theorem 5.6 is satisfied in the case where $T\coloneqq F,\,Y\coloneqq\widehat{\mathbb{C}}\smallsetminus F^{-1}(\operatorname{post}{F}),\,J\coloneqq\deg{F}\cdot\mathbbm{1}_{\widehat{\mathbb{C}}}$, and $\phi\coloneqq\phi_{0}$. Moreover, since $F$ is an expanding Thurston map, we have $\operatorname{card}\mathopen{\big(}F^{-1}(x)\smallsetminus F^{-1}(\operatorname{post}{F})\bigr)\mathclose{\big)}=\deg{F}$ for each $x\in F\bigl(\widehat{\mathbb{C}}\smallsetminus F^{-1}(\operatorname{post}{F})\bigr)$. Hence, property (ii) in Theorem 5.6 is satisfied in the case where $T\coloneqq F,\,Y\coloneqq\widehat{\mathbb{C}}\smallsetminus F^{-1}(\operatorname{post}{F})$, and $J\coloneqq\deg{F}\cdot\mathbbm{1}_{\widehat{\mathbb{C}}}$. This verifies Assumption (iii).
@@
Assumption (v): Now we demonstrate that $\mathrm{Per}(F)\cap\operatorname{post}{F}$ is the set of periodic points of $F$ in $F^{-1}(\operatorname{post}{F})$, namely, $\mathrm{Per}(F)\cap\operatorname{post}{F}=\mathrm{Per}(F)\cap F^{-1}(\operatorname{post}{F})$. Indeed, this immediately follows from the property that $F(x)=x$ for each $x\in\mathrm{Per}(F)\cap\operatorname{post}{F}$. Since $\{y_{i}\}_{i\in I}$ is uniformly computable, $\{p_{n,i}\}_{(n,i)\in H}$, with $p_{n,i}=y_{i}$ for all $(n,i)\in H$, is also uniformly computable. Moreover, by the definition of $\{y_{i}\}_{i=1}^{N}$, we have that $\{p_{n,i}:(n,i)\in H_{n}\}=\{y_{i}:i\in I\}=\mathrm{Per}(F)\cap\operatorname{post}{F}$. Therefore, Claim 2 follows.
@@
Proof of Claim 3. Recall that $\{\mathbb{B}_{i}\}_{i=1}^{N}$ is uniformly lower semi-computable open and $F$ is a computable function. Define $U_{n,i,1}\coloneqq\mathbb{B}_{i}$ and $U_{n,i,k+1}\coloneqq F^{-1}(U_{n,i,k})\cap\mathbb{B}_{i}$ by induction for each $n,\,k\in\mathbb{N}$, and each $i\in I$. Then by Corollary 3.22, we have $\{U_{n,i,k}\}_{((n,i),k)\in H\times\mathbb{N}}$ is uniformly lower semi-computable open. Recall that $y_{i}$ is a fixed point of $F$ and $y_{i}\in\mathbb{B}_{i}$ for each $i\in I$. Then we obtain that $y_{i}\in U_{n,i,k}$ for all $n,\,k\in\mathbb{N}$, and $i\in I$.
@@
Next, we set $t_{n,i,k}=\bigl(\deg_{F}(y_{i})\big/\deg{F}\bigr)^{k-1}$ for each $n,\,k\in\mathbb{N}$, and each $i\in I$. Then by [Li17, Lemma 4.27], $\deg_{F}(y_{i})<\deg{F}$ for each $i\in I$. Thus, for all $n\in\mathbb{N}$ and $i\in I$, we obtain that $\inf_{k\in\mathbb{N}}t_{n,i,k}=\inf_{k\in\mathbb{N}}\bigl(\deg_{F}(y_{i})\big/\deg{F}\bigr)^{k}=0$. Moreover, since $\deg_{F}(y_{i})\big/\deg{F}\in\mathbb{Q}$ for each $i\in I$, we have that $\{t_{n,i,k}\}_{((n,i),k)\in H\times\mathbb{N}}$ is a sequence of uniformly computable real numbers.
@@
Finally, we consider an arbitrary pair of $n\in\mathbb{N}$ and $i\in I$ and prove that $\mu_{0}(U_{n,i,k})\leqslant t_{n,i,k}$ for each $k\in\mathbb{N}$. Recall that $y_{i}\in\operatorname{post}{F}$ and $F(y_{i})=y_{i}$. For each $m\in\mathbb{N}$, we define the set $\mathbf{\mathfrak{X}}^{m}(F,\mathcal{C},y_{i})$ of $m$-tiles of $F$ at $y_{i}$ as $\mathbf{\mathfrak{X}}^{m}(F,\mathcal{C},y_{i})\coloneq\{X\in\mathbf{X}^{m}(F,\mathcal{C}):y_{i}\in X\}$, then by Remark 6.10, we obtain that $\overline{W}^{m}(y_{i})=\bigcup\mathbf{\mathfrak{X}}^{m}(F,\mathcal{C},y_{i})$ and $\operatorname{card}\mathopen{(}\mathbf{\mathfrak{X}}^{m}(F,\mathcal{C},y_{i})\mathclose{)}=2\mathopen{\big(}\deg_{F}(y_{i})\mathclose{\big)}^{m}$. In particular, (6.6) implies that $F\bigl(W^{m+1}(y_{i})\bigr)=W^{m}(y_{i})$ for each $m\in\mathbb{N}$. Employing [BM17, Proposition 17.12] (recall that the Jordan curve $\mathcal{C}$ is $F$-invariant) and the fact that $\mu_{0}(W^{m}(y_{i}))=\mu_{0}\bigl(\overline{W}^{m}(y_{i})\bigr)$ for each $m\in\mathbb{N}$ (cf. [Li18, Proposition 7.1]), we deduce that $\mu_{0}\bigl(W^{m+1}(y_{i})\bigr)=\mu_{0}(W^{m}(y_{i}))\cdot\bigl(\deg_{F}(y_{i})\big/\deg{F}\bigr)$ for each $m\in\mathbb{N}$. Then it follows from the induction that $\mu_{0}\bigl(W^{m+1}(y_{i})\bigr)=\bigl(\deg_{F}(y_{i})\big/\deg{F}\bigr)^{m}\mu_{0}\bigl(W^{1}(y_{i})\bigr)\leqslant\bigl(\deg_{F}(y_{i})\big/\deg{F}\bigr)^{m}$ for each $m\in\mathbb{N}$. Since $\mathbb{B}_{i}\subseteq W^{1}(y_{i})$, by the definition of $\{U_{n,i,k}\}_{k\in\mathbb{N}}$, $U_{n,i,k}\subseteq W^{k}(y_{i})$ for each $k\in\mathbb{N}$. Hence, $\mu_{0}(U_{n,i,k})\leqslant\mu_{0}\bigl(W^{k}(y_{i})\bigr)\leqslant\bigl(\deg_{F}(y_{i})\big/\deg{F}\bigr)^{k-1}=t_{n,i,k}$ for each $k\in\mathbb{N}$.
Therefore, we have shown Claim 3.
@@
Now we turn to show the original statement. Indeed, by Claims 1, 2, and 3, it follows from Theorem 5.9 that $\{\mu_{n}\}_{n\in\mathbb{N}}$, with $\mu_{n}=\mu_{0}$ for all $n\in\mathbb{N}$, is a sequence of uniformly computable measures. Thus, $\mu_{0}$ is a computable measure. ∎
@@
As concrete applications of Theorem 1.4, we analyze two expanding Thurston maps $g_{1}$ and $g_{2}$ with periodic critical points. We verify that both maps satisfy the hypotheses of Theorem 1.4, thereby establishing the computability of their measures of maximal entropy $\mu_{1}$ and $\mu_{2}$ (Corollary 6.11). Crucially, since both maps have periodic critical points, their measure-theoretic entropy functions fail to be upper semi-continuous (Theorem 6.4), precluding a direct application of Theorem 1.1.
@@
Our first example $g_{1}$ is derived from the barycentric subdivision rule as defined in [BM17, Example 12.21]. We construct a sphere $S_{\triangle}$ by gluing two equilateral (Euclidean) triangles along their boundaries (see Figure 6.1). These two triangles serve as 0-tiles, with their shared boundary forming a $g_{1}$-invariant Jordan curve containing the postcritical set $\{A,\,B,\,C\}$ of $g_{1}$. Subdividing each 0-tile via bisectors yields six smaller triangles, producing twelve $1$-tiles. The map $g_{1}$ sends each $1$-tile linearly onto a 0-tile, and has a fixed critical point. Recursively, $n$-tiles are generated by analogously subdividing (n1)(n-1)-tiles via bisectors, and are Euclidean triangles. It is not difficult to see that the diameters of $n$-tiles decay to zero as $n$ tends to $\infty$, confirming that $g_{1}$ is an expanding Thurston map.
@@
To further demonstrate the scope of Theorem 1.4 while illustrating the complexity of expanding Thurston maps, we construct a more intricate example $g_{2}$. The map $g_{2}$ acts on the same polyhedral sphere $S_{\triangle}$ with the postcritical set $\{A,\,B,\,C\}$. While $g_{1}$ subdivides each 0-tile into six congruent smaller triangles, $g_{2}$ implements a finer subdivision rule: each 0-tile splits into eight triangles following the structure illustrated in Figure 6.2. The map $g_{2}$ linearly maps each $1$-tile onto a 0-tile and has a fixed critical point. Recursive iteration of this subdivision generates $n$-tiles whose diameters decay to zero as $n$ tends to $\infty$, showing that $g_{2}$ is an expanding Thurston map.

### v2

@@ 6.3. Computable expanding Thurston maps
We begin by constructing a recursively enumerable set $K$ and a sequence $\{Y_{n,k}\}_{(n,k)\in L}$ of uniformly lower semi-computable open sets, where $L=\mathbb{N}\times K$. Now we write $\mathcal{S}\bigl(\widehat{\mathbb{C}}\bigr)=\{s_{i}\}_{i\in\mathbb{N}}$. Since $\operatorname{post}{F}$ is a finite set of computable points, by Proposition 3.20 (i), (ii), and (v), $\widehat{\mathbb{C}}\smallsetminus\operatorname{post}{F}$ is a nonempty lower semi-computable open set. By Proposition 3.8, there exists a recursively enumerable set $I\subseteq\mathbb{N}$ such that $\mathbb{Q}\bigl(\widehat{\mathbb{C}}\bigr)\smallsetminus\operatorname{post}{F}=\{s_{i}:i\in I\}$. We define $r_{i}\coloneqq\sigma(s_{i},\operatorname{post}{F})>0$ for each $i\in I$. Thus, by Definition 3.3, $\{r_{i}\}_{i\in I}$ is uniformly computable. Consequently, by Proposition 3.10, $\{D_{i}\}_{i\in I}$ is uniformly lower semi-computable open, where $D_{i}\coloneqq B(s_{i},r_{i})$ for each $i\in I$. As $F$ is computable, $\bigl\{F^{-1}(D_{i})\bigr\}_{i\in I}$ is also uniformly lower semi-computable. By Lemma 6.9, there exists a recursively enumerable set $K\subseteq\mathbb{N}\times I$ and a sequence $\{V_{k}\}_{k\in K}$ of uniformly lower semi-computable open sets that are connected such that $\bigcup_{k\in K_{i}}V_{k}=F^{-1}(D_{i})$ for each $i\in I$.
@@
In order to show that for each pair of distinct integers $1\leqslant l<l^{\prime}\leqslant b$,
$h_{l}(D_{i})\cap h_{l^{\prime}}(D_{i})=\emptyset,$ (6.5)
we argue by contradiction and assume that (6.5) is not true for some pair of distinct integers $1\leqslant l<l^{\prime}\leqslant b$. Then there exist $z,\,z^{\prime}\in D_{i}$ with $h_{l}(z)=h_{l^{\prime}}(z^{\prime})$. Hence, $z=F(h_{l}(z))=F(h_{l^{\prime}}(z^{\prime}))=z^{\prime}$. Combined with [BM17, Lemma A.6 (i)], this implies that $h_{l}=h_{l^{\prime}}$, which contradicts $l<l^{\prime}$. Thus, (6.5) is true for each pair of distinct integers $1\leqslant l<l^{\prime}\leqslant b$.
@@
We can enumerate $\mathrm{Per}(F)\cap\operatorname{post}{F}$ as a sequence $\{y_{i}\}_{i=1}^{N}$ of uniformly computable (mutually distinct) points, since it is a finite set of computable points.
We now construct a sequence $\{\mathbb{B}_{i}\}_{i=1}^{N}$ of uniformly lower semi-computable open sets satisfying that $\{y_{i}\}\subseteq\mathbb{B}_{i}\subseteq\overline{W}^{1}(y_{i})$ for each integer $1\leqslant i\leqslant N$. Fix an arbitrary integer $1\leqslant i\leqslant N$. By [BM17, Lemma 5.28 (i)], there exists $R_{i}\in\mathbb{Q}^{+}$ such that $B(y_{i},R_{i})\subseteq\overline{W}^{1}(y_{i})$. Since $y_{i}$ is a computable point, by Proposition 3.10, $\mathbb{B}_{i}\coloneqq B(y_{i},R_{i})$ is a lower semi-computable open set with $\{y_{i}\}\subseteq\mathbb{B}_{i}\subseteq\overline{W}^{1}(y_{i})$. Hence, by Definition 3.7, $\{\mathbb{B}_{i}\}_{i=1}^{N}$ is uniformly lower semi-computable open.
@@
Proof of Claim 2. Now we verify items (ii) to (v) in the Assumptions in Subsection 5.3.
Item (ii): Since $\phi_{0}$ is a constant function, this assumption holds.
@@
Item (iv): Note that $\widehat{\mathbb{C}}$ is recursively compact in $(X,\,\rho,\,\mathcal{S})$ by Proposition 6.5. Hence, by definitions, it suffices to show that $\operatorname{card}\mathopen{\big(}F^{-1}(\operatorname{post}{F})\mathclose{\big)}<+\infty$ and $\mathcal{E}_{0}\bigl(F,\phi_{0};\widehat{\mathbb{C}}\smallsetminus F^{-1}(\operatorname{post}{F})\bigr)\cap\mathcal{P}\bigl(X;\widehat{\mathbb{C}}\smallsetminus F^{-1}(\operatorname{post}{F})\bigr)=\{\mu_{0}\}$. By Theorem 6.3 and [BM17, Corollary 17.3], we have $\mathcal{E}(F,\phi_{0})=\mathcal{E}(f,\phi_{0})=\{\mu_{0}\}$. Moreover, it follows from [Li18, Corollary 7.4] that $\mu_{0}\bigl(F^{-1}(\operatorname{post}{F})\bigr)=\mu_{0}(\operatorname{post}{F})=0$. Combined with (19.5) in [BM17] and (5.16), this implies that $\mathcal{E}_{0}\bigl(F,\phi_{0};\widehat{\mathbb{C}}\smallsetminus F^{-1}(\operatorname{post}{F})\bigr)=\{\mu_{0}\}$.
@@
Item (iii): Since $F$ is an expanding Thurston map, by [BM17, Corollary 17.2], we have $h_{\mathrm{top}}(F)=\log(\deg{F})$. Suppose $h=\phi_{0}$. Then property (i) in Theorem 5.6 is satisfied in the case where $T\coloneqq F,\,Y\coloneqq\widehat{\mathbb{C}}\smallsetminus F^{-1}(\operatorname{post}{F}),\,J\coloneqq\deg{F}\cdot\mathbbm{1}_{\widehat{\mathbb{C}}}$, and $\phi\coloneqq\phi_{0}$. Moreover, since $F$ is an expanding Thurston map, we have $\operatorname{card}\mathopen{\big(}F^{-1}(x)\smallsetminus F^{-1}(\operatorname{post}{F})\mathclose{\big)}=\deg{F}$ for each $x\in F\bigl(\widehat{\mathbb{C}}\smallsetminus F^{-1}(\operatorname{post}{F})\bigr)$. Hence, property (ii) in Theorem 5.6 is satisfied in the case where $T\coloneqq F,\,Y\coloneqq\widehat{\mathbb{C}}\smallsetminus F^{-1}(\operatorname{post}{F})$, and $J\coloneqq\deg{F}\cdot\mathbbm{1}_{\widehat{\mathbb{C}}}$. This verifies Item (iii) in the Assumptions.
@@
Item (v): Now we demonstrate that $\mathrm{Per}(F)\cap\operatorname{post}{F}$ is the set of periodic points of $F$ in $F^{-1}(\operatorname{post}{F})$, namely, $\mathrm{Per}(F)\cap\operatorname{post}{F}=\mathrm{Per}(F)\cap F^{-1}(\operatorname{post}{F})$. Indeed, this immediately follows from the property that $F(x)=x$ for each $x\in\mathrm{Per}(F)\cap\operatorname{post}{F}$. Since $\{y_{i}\}_{i\in I}$ is uniformly computable, $\{p_{n,i}\}_{(n,i)\in H}$, with $p_{n,i}=y_{i}$ for all $(n,i)\in H$, is also uniformly computable. Moreover, by the definition of $\{y_{i}\}_{i=1}^{N}$, we have that $\{p_{n,i}:(n,i)\in H_{n}\}=\{y_{i}:i\in I\}=\mathrm{Per}(F)\cap\operatorname{post}{F}$. Therefore, Claim 2 follows.
@@
Proof of Claim 3. Recall that $\{\mathbb{B}_{i}\}_{i=1}^{N}$ is uniformly lower semi-computable open and $F$ is a computable function. Define $U_{n,i,1}\coloneqq\mathbb{B}_{i}$ and $U_{n,i,k+1}\coloneqq F^{-1}(U_{n,i,k})\cap\mathbb{B}_{i}$ recursively for each $n,\,k\in\mathbb{N}$, and each $i\in I$. Then by Corollary 3.22, we have $\{U_{n,i,k}\}_{((n,i),k)\in H\times\mathbb{N}}$ is uniformly lower semi-computable open. Recall that $y_{i}$ is a fixed point of $F$ and $y_{i}\in\mathbb{B}_{i}$ for each $i\in I$. Then we obtain that $y_{i}\in U_{n,i,k}$ for all $n,\,k\in\mathbb{N}$, and $i\in I$.
@@
Next, we set $t_{n,i,k}\coloneqq(\deg_{F}(y_{i})/\deg{F})^{k-1}$ for each $n,\,k\in\mathbb{N}$, and each $i\in I$. Then by [Li17, Lemma 4.27], $\deg_{F}(y_{i})<\deg{F}$ for each $i\in I$. Thus, for all $n\in\mathbb{N}$ and $i\in I$, we obtain that $\inf_{k\in\mathbb{N}}t_{n,i,k}=\inf_{k\in\mathbb{N}}(\deg_{F}(y_{i})/\deg{F})^{k}=0$. Moreover, since $\deg_{F}(y_{i})/\deg{F}\in\mathbb{Q}$ for each $i\in I$, we have that $\{t_{n,i,k}\}_{((n,i),k)\in H\times\mathbb{N}}$ is a sequence of uniformly computable real numbers.
@@
Finally, we consider an arbitrary pair of $n\in\mathbb{N}$ and $i\in I$ and prove that $\mu_{0}(U_{n,i,k})\leqslant t_{n,i,k}$ for each $k\in\mathbb{N}$. Recall that $y_{i}\in\operatorname{post}{F}$ and $F(y_{i})=y_{i}$. For each $m\in\mathbb{N}$, it follows from [BM17, Lemma 5.28 (ii)] that $\operatorname{card}\mathopen{\big(}\overline{W}^{m}(y_{i})\mathclose{\big)}=2(\deg_{F}(y_{i}))^{m}$. In particular, [BM17, Proposition 5.16] implies that $F\bigl(\overline{W}^{m+1}(y_{i})\bigr)=\overline{W}^{m}(y_{i})$ for each $m\in\mathbb{N}$. Employing [BM17, Proposition 17.12] and [Li18, Proposition 7.1] (recall that the Jordan curve $\mathcal{C}$ is $F$-invariant), we deduce that $\mu_{0}\bigl(\overline{W}^{m+1}(y_{i})\bigr)=\mu_{0}\bigl(\overline{W}^{m}(y_{i})\bigr)\cdot(\deg_{F}(y_{i})/\deg{F})$ for each $m\in\mathbb{N}$. Then it follows from the induction that $\mu_{0}\bigl(\overline{W}^{m+1}(y_{i})\bigr)=(\deg_{F}(y_{i})/\deg{F})^{m}\mu_{0}\bigl(\overline{W}^{1}(y_{i})\bigr)\leqslant(\deg_{F}(y_{i})/\deg{F})^{m}$ for each $m\in\mathbb{N}$. Since $\mathbb{B}_{i}\subseteq\overline{W}^{1}(y_{i})$, by the definition of $\{U_{n,i,k}\}_{k\in\mathbb{N}}$, $U_{n,i,k}\subseteq\overline{W}^{k}(y_{i})$ for each $k\in\mathbb{N}$. Hence, $\mu_{0}(U_{n,i,k})\leqslant\mu_{0}\bigl(\overline{W}^{k}(y_{i})\bigr)\leqslant(\deg_{F}(y_{i})/\deg{F})^{k-1}=t_{n,i,k}$ for each $k\in\mathbb{N}$. Therefore, we have shown Claim 3.
@@
We are now in a position to complete the proof of Theorem 1.4. By Claims 1, 2, and 3, it follows from Theorem 5.10 that the constant sequence $\{\mu_{n}\}_{n\in\mathbb{N}}$, defined by $\mu_{n}=\mu_{0}$ for all $n\in\mathbb{N}$, is a sequence of uniformly computable measures. Thus, $\mu_{0}$ is computable. ∎
@@
As concrete applications of Theorem 1.4, we analyze two expanding Thurston maps $g_{1}$ and $g_{2}$ with periodic critical points. We verify that both maps satisfy the hypotheses of Theorem 1.4, thereby establishing the computability of their measures of maximal entropy $\mu_{1}$ and $\mu_{2}$ (Corollary 6.10). Crucially, since both maps have periodic critical points, their measure-theoretic entropy functions fail to be upper semicontinuous (Theorem 6.4), precluding a direct application of Theorem 1.1.
@@
Our first example $g_{1}$ is derived from the barycentric subdivision rule as defined in [BM17, Example 12.21]. We construct a sphere $S_{\triangle}$ by gluing two equilateral (Euclidean) triangles along their boundaries (see Figure 6.1). These two triangles serve as 0-tiles. Subdividing each 0-tile via bisectors yields six smaller triangles, producing twelve $1$-tiles. Recursively, $n$-tiles are generated by analogously subdividing (n1)(n-1)-tiles via bisectors, and are Euclidean triangles. The map $g_{1}\colon S_{\triangle}\rightarrow S_{\triangle}$ is defined as a piecewise linear map on $S_{\triangle}$ in the following way: the orientation-preserving branched covering map $g_{1}$ is affine on each $1$-tile and maps the $1$-tile linearly onto a 0-tile in a way such that $g_{1}$ is continuous on $S_{\triangle}$. Then $g_{1}$ has a fixed critical point, and the shared boundary of 0-tiles forming a $g_{1}$-invariant Jordan curve containing the postcritical set $\{A,\,B,\,C\}$ of $g_{1}$. It is not difficult to see that the diameters of $n$-tiles decay to zero as $n$ tends to $\infty$, confirming that $g_{1}$ is an expanding Thurston map.
@@
To further demonstrate the scope of Theorem 1.4 while illustrating the complexity of expanding Thurston maps, we construct a more intricate example $g_{2}$. The map $g_{2}$ acts on the same polyhedral sphere $S_{\triangle}$ with the postcritical set $\{A,\,B,\,C\}$. While $g_{1}$ subdivides each 0-tile into six congruent smaller triangles, $g_{2}$ implements a finer subdivision rule: each 0-tile splits into eight triangles following the structure illustrated in Figure 6.2. Similarly, the orientation-preserving branched covering map $g_{2}\colon S_{\triangle}\rightarrow S_{\triangle}$ is defined as a piecewise linear map on $S_{\triangle}$ that maps each $1$-tile onto a 0-tile linearly. The map $g_{2}$ also has a fixed critical point. Recursive iteration of this subdivision generates $n$-tiles whose diameters decrease to zero as $n$ tends to $\infty$, showing that $g_{2}$ is an expanding Thurston map.